OpenStudy (anonymous):

Solve. 64x = 16x – 1 a. x = –2 b. x = –1 c. x = -1/4 d. x = -1/3

OpenStudy (anonymous):

64x = 16x -1 48x = -1 x=-1/48 Are you sure the choices or problem is correct?

OpenStudy (anonymous):

OH! Sorry. 64^x=16^(x-1)

OpenStudy (anonymous):

AH! Ok to solve for a variable in the exponent, you want to change both sides to having the same base. So you want a common factor between 64 and 16. Which would be? :)

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

You could use 4; you could also use 8. We'll do it your way :) So 4 to what power gets 64? and 4 to what power gets 16?

OpenStudy (anonymous):

64/4 is 16 and 16/4 is 4.

OpenStudy (anonymous):

4^3 = 64. I was meaning the exponent; not division :) SO it would look like this 4^3x = 4^(2x-2) Then you can set 3x=2x-2 and solve for x=-2 :) does that make sense?

OpenStudy (anonymous):

Yep! Thanks!

OpenStudy (anonymous):

Great :) anytime!